# MMWorld: Towards Multi-discipline Multi-faceted World Model Evaluation in Videos

### v1

@@ 1 Introduction
  • We observe that existing MLLMs still face substantial challenges posed by MMWorld. Even the best performer, GPT-4V, can only achieve a 52.30% overall accuracy, and four MLLMs particularly trained on videos perform worse than random chance.
@@ 5 Conclusion
Appendix A Overview
We host the project website on https://mmworld-bench.github.io/. The benchmark and code implementations can be found at https://github.com/eric-ai-lab/MMWorld. The link to Croissant metadata record documenting the dataset/benchmark available for viewing and downloading is available at https://github.com/eric-ai-lab/MMWorld/blob/main/data/croissant_data.json. This Appendix is organized as follows:
@@ Appendix B Additional Results
B.1 Results for the Two Different Evaluation Strategies
In Table 7, we give additional evaluation results for different methods evaluated in this paper. For closed-source models, one evaluation framework is the one used in the main paper, which involves utilizing GPT-4V as a judger. The process consists of presenting GPT-4V with the question, a corresponding answer generated by the baseline model, and the set of possible options. GPT-4V then assesses whether the model-generated answer is accurate within the given context. This approach ensures consistency across our evaluations and allows for a direct comparison of the baseline models’ performance against GPT-4V’s judgement; Another is open-ended generation where we employ a two-step methodology. We first prompt each model to do open-ended generation. Subsequently, we prompt the model to align its generative response with one of the predefined options: ‘a’, ‘b’, ‘c’, or ‘d’.
B.2 Results Across Different Seed for Each Model
@@
B.3 Results from Amazon Turkers
@@
Appendix C Implementation Details
We use the optimum number of video frames and report the performance in the main paper. The numbers of the sampled frames are 10 for GPT-4V/o, 8 for Video-LLaVA, 16 for VideoChat2, 64 for ST-LLM, 32 for Chat-UniVi-V1.5, 32 for LLaVA-NeXT-Video, 4 for Qwen-VL-Chat/Max, and 10 for InternVL-Chat-V1.5. For VideoChat, we set ‘max_frames’ to 100.
For closed-source models, for both Gemini Pro and GPT-4V, we use the default settings provided by their official APIs. We use Katna 333 https://github.com/keplerlab/katna to extract key video frames as input to these two models. The Gemini Pro is set to process visual input and configured with safety settings to filter a range of harmful content. The configuration thresholds are set to ‘BLOCK_NONE’. For PandaGPT, we set ‘top_p’ to 0.7, and ‘temperature’ to 0.5. For VideoChat, we set ‘max_frames’ to 100. For LWM, we use the LWM-Chat-1M variant. For X-Instruct-BLIP, the model is implemented using four image frames. For Otter, we use the video variant. We use GPT-4-32K as the judge for judging whether the model answer is correct when it can not mapped to the option letter using the rule-based method. The prompt provided to GPT-4-32K is structured as follows: "I will present a response from a question-answering model alongside several answer options. Your task is to evaluate the response and determine which of the following options it most closely aligns with, denoting the most similar option by its corresponding letter (a, b, c, or d).".
Query Generation in Synthetic Data Generation Pipeline
Table 7: Performance of various baselines across different disciplines.
@@ D.1 Quality of Data
We also hired students from campus to do human evaluation on subset of the data. The results are shown in Table 9. The performance of the human evaluators did not surpass that of GPT-4V and Gemini-Pro. This outcome underscores the challenging nature of the dataset, which often necessitates specialized domain knowledge that our evaluators—primarily non-experts—found demanding. These results highlight the complexity of the questions and the potential necessity for domain-specific understanding to achieve high accuracy
@@ D.2 Quality of Using GPT as the Judger
For a comprehensive assessment of GPT-4V’s accuracy when using it as the judger, we devised a human evaluation protocol also resort to Amazon Mechanical Turk, as visualized in Figure 8. The evaluators present a series of statements derived from the video, and GPT-4V is tasked with selecting the most accurate answer from a set of multiple-choice questions. The interface is designed to mimic the complexity of real-world scenarios where GPT-4V must infer, deduce, or predict outcomes based on its understanding of the content. Each section comprises a targeted question that prompts GPT-4V to consider the video from various analytical perspectives, such as character motivations, narrative developments, and broader thematic implications. Through this interface, evaluators can efficiently gauge GPT-4V’s performance across different types of questions—ranging from those requiring surface-level understanding to ones necessitating deeper analytical thought. This protocol enables us to quantify GPT-4V’s proficiency in tasks that traditionally rely on human cognitive abilities, offering insights into the model’s understanding of nuanced and context-dependent information.
@@ Appendix E Error Analysis
We show in Figure 15161718 some error cases of Question Understanding Error, Audio Understanding Error, Visual Perception Error, Hallucinations, Reasoning Error, Lack of Domain Knowledge, and Reject to Answer respectively from methods evaluated on MMWorld.
@@ Appendix G Additional Data Statistics
For human annotated dataset, the length of each video was capped at approximately two minutes. The statistical distribution of the disciplines within the dataset is as follows:

### v3

@@ 1 Introduction
  • We observe that existing MLLMs still face substantial challenges posed by MMWorld. Even the best performer, GPT-4o, can only achieve a 52.30% overall accuracy, and four MLLMs particularly trained on videos perform worse than random chance.
@@ 5 Conclusion
Appendix A Overview of the Appendix
We host the project website on https://mmworld-bench.github.io/. The benchmark and code implementations can be found at https://github.com/eric-ai-lab/MMWorld. The link to Croissant metadata record documenting the dataset/benchmark available for viewing and downloading is available at https://github.com/eric-ai-lab/MMWorld/blob/main/data/croissanta_hf_data.json. This Appendix is organized as follows:
@@ Appendix B Additional Results
B.1 Results Across Different Seed for Each Model
@@
GPT-4o-seed 1 [OpenAI, 2024] 47.10 92.31 75.11 81.25 65.03 72.34 78.31 62.22
GPT-4o-seed 2 [OpenAI, 2024] 46.58 90.91 69.78 84.38 65.73 75.53 83.13 61.77
GPT-4o-seed 3 [OpenAI, 2024] 49.94 90.21 76.44 84.38 58.04 78.72 79.52 63.63
Claude-3.5-seed 1 [Anthropic, 2024] 54.32 64.34 59.11 53.12 30.77 59.57 59.04 54.27
Claude-3.5-seed 2 [Anthropic, 2024] 54.32 63.64 61.33 54.17 30.77 58.51 59.04 54.52
Claude-3.5-seed 3 [Anthropic, 2024] 55.10 63.64 59.11 56.25 31.47 58.51 60.24 54.84
@@
B.2 Results from Amazon Turkers
@@
B.3 Results for the Two Different Evaluation Strategies
Table 7: Performance of different MLLMs across different disciplines.
@@ B.3 Results for the Two Different Evaluation Strategies
In Table 7, we give additional evaluation results for different MLLMs evaluated in this paper. For closed-source models, the evaluation pipeline is the one used in the main paper, which involves utilizing GPT-4V as a judger. The process consists of presenting GPT-4V with the question, a corresponding answer generated by the baseline model, and the set of possible options. GPT-4V then assesses whether the model-generated answer is accurate within the given context; Another is open-ended generation where we employ a two-step methodology. We first prompt each model to do open-ended generation. Subsequently, we prompt the model to align its generative response with one of the predefined options: ‘a’, ‘b’, ‘c’, or ‘d’.
B.4 Detailed Results on Multi-faceted Reasoning
Table 8: Detailed results of different MLLMs on multi-faceted reasoning.
Model Explanation Counterfactual Future Domain Attribution Temporal
Thinking Prediction Expertise Understanding Understanding
Proprietary Models
GPT-4V 44.90 64.90 78.59 61.07 59.61 27.17
Gemini Pro 48.58 65.49 65.45 53.87 43.92 24.65
Open-source Models
Video-LLaVA-7B 42.46 42.55 64.96 47.86 36.86 34.45
VideoChat-7B 41.66 43.73 45.74 40.95 30.59 25.77
ImageBind-LLM-7B 29.51 26.86 50.61 33.93 34.90 19.89
PandaGPT-7B 29.55 37.45 46.47 33.93 26.27 28.01
ChatUnivi-7B 33.91 48.82 61.80 45.95 33.33 22.97
VideoLLaMA-2-13B 10.55 23.92 25.30 16.31 8.63 6.16
X-Instruct-BLIP-7B 23.05 15.29 27.25 21.07 24.31 11.20
LWM-1M-JAX 11.62 18.82 30.66 17.98 21.57 7.00
Otter-7B 16.91 10.98 15.82 13.10 17.65 9.52
mPLUG-Owl-7B 35.20 49.61 55.47 47.74 24.71 20.17
In Table 8, we give detailed performance numbers of different MLLMs on multi-faceted reasoning corresponding to Figure 4 in the main paper.
Appendix C Implementation Details
We use the optimum number of video frames and report the performance in the main paper. The numbers of the sampled frames are 10 for GPT-4V/o and Gemini Pro, 8 for Video-LLaVA, 32 for ChatUniVi. For closed-source models, for both Gemini Pro and GPT-4V, we use the default settings provided by their official APIs. We use Katna 333 https://github.com/keplerlab/katna to extract key video frames as input to these two models. The Gemini Pro is set to process visual input and configured with safety settings to filter a range of harmful content. The configuration thresholds are set to ‘BLOCK_NONE’. For PandaGPT, we set ‘top_p’ to 0.7, and ‘temperature’ to 0.5. For VideoChat, we set ‘max_frames’ to 100. For LWM, we use the LWM-Chat-1M variant. For X-Instruct-BLIP, the model is implemented using four image frames. For Otter, we use the video variant. We use GPT-4-32K as the judge for judging whether the model answer is correct when it can not mapped to the option letter using the rule-based method. The prompt provided to GPT-4-32K is structured as follows: "I will present a response from a question-answering model alongside several answer options. Your task is to evaluate the response and determine which of the following options it most closely aligns with, denoting the most similar option by its corresponding letter (a, b, c, or d).".
Query Generation in Synthetic Data Generation Pipeline
@@ D.1 Quality of Data
We also hired students from campus to do human evaluation on subset of the data. The results are shown in Table 10. The performance of the human evaluators did not surpass that of GPT-4V and Gemini-Pro. This outcome underscores the challenging nature of the dataset, which often necessitates specialized domain knowledge that our evaluators—primarily non-experts—found demanding. These results highlight the complexity of the questions and the potential necessity for discipline-specific understanding to achieve high accuracy
@@ D.2 Quality of Using GPT as the Judger
For a comprehensive assessment of GPT-4V’s accuracy when using it as the judger, we devised a human evaluation protocol also resort to Amazon Mechanical Turk, as visualized in Figure 8. The evaluators present a series of statements derived from the video, and GPT-4V is tasked with selecting the most accurate answer from a set of multiple-choice questions. Through this interface, human evaluators can efficiently gauge GPT-4V’s performance across different types of questions—when using it as the judger.
@@ Appendix E Error Analysis
We show in Figure 15161718 some error cases of Question Understanding Error, Audio Understanding Error, Visual Perception Error, Hallucinations, Reasoning Error, Lack of Domain Knowledge, and Reject to Answer respectively from MLLMs evaluated on MMWorld.
@@ Appendix G Additional Data Statistics
For human annotated dataset, the length of each video was capped at approximately two minutes. The statistical distribution of the disciplines within the dataset for this part is as follows: